OpenStudy (anonymous):

if you have a triangle with all 3 sides measuring 4 cm then what would be the measure of all 3 angles??

OpenStudy (anonymous):

If all of the sides are the same length then you have an equilateral triangle, which means all of the angles are equal as well. Every triangle's angles add up to 180 degrees and so you just divide 180 by 3 to get 60 degrees for each angle on an equilateral triangle.

OpenStudy (anonymous):

friend you can use cosine law.

OpenStudy (anonymous):

cosA = (-a^2+b^2+c^2)/(2bc) cosB = (a^2-b^2+c^2)/(2ac) cosC = (a^2+b^2-c^2)/(2ab).

OpenStudy (anonymous):

or i could just take sanguines advice:D

OpenStudy (anonymous):

my method is easy.

OpenStudy (anonymous):

divide 180 by 3 is not easy lol his was easier i have done cos tan and sine and they are not very fast methods:D

OpenStudy (anonymous):

its easy 180/3=60

OpenStudy (anonymous):

that is what the first guy said?!?

OpenStudy (anonymous):

did you try my method??

OpenStudy (anonymous):

180/3=60